A police car of mass 2.4x103kg is travelling on the highway when the officers receive an emergency call. They increase the speed of the car to 33 m/s.  The increase in speed results in 3.1 x 10 5 J of work done on the car. Determine the initial speed of the police car in kilometers per hour?


1
Expert's answer
2022-01-17T09:59:49-0500

Given:

"m=\rm2.4*10^3\: kg"

"v_f=\rm 33\: m/s"

"W=3.1*10^5\:\rm J"


The energy-work theorem says

"W=\Delta E=\frac{mv_f^2}{2}-\frac{mv_i^2}{2}"

Hence, the initial speed

"v_i=\sqrt{v_f^2-2W/m}"

"v_i=\sqrt{33^2-2*3.1*10^5/2.4*10^3}=29\:\rm m/s"